Water is boiling in an aluminum pan placed on an electrical element on a stovetop. The sauce pan has a bottom that is 0.800 cm thick and 14.0 cm in diameter. The boiling water is evaporating at the rate of 1.00 g/s. What is the temperature difference across (through) the bottom of the pan?


1
Expert's answer
2021-05-27T10:21:34-0400

Gives

Thickness(d)=0.800Cm "=8\times10^{-3}m"


Diameter=14.0 cm

Radius="7\times10^{-2}m"

"\frac{m}{t}=1.0\times10^{-3}kg/sec"

KAl=220J/sec.m°c

L=2256"\times10^{3}J/kg"


"\frac{dQ}{dt}=\frac {kA({T_2-T_1)}}{d}"

Q=mL

"T_2-T_1=\frac{\frac{dQ}{dt}\times d}{kA}"

"T_2-T_1=\frac{L\frac{dm}{dt}\times d}{kA}\rightarrow(1)"

Put Value

"T_2-T_1=\frac{2256\times10^3\times\frac{10^{-3}}{1}\times 8\times10^{-3}}{220\times3.14\times(7\times10^{-2})^2}"

"T_2-T_1=5.33\u00b0C"